OpenStudy (nick5006):

Given f(x)=1/"absolute value x" with the following: f(2+h) where h>0.

OpenStudy (nick5006):

\[f(x)=1/\left| x \right|\]

OpenStudy (nick5006):

I think the answer is f(x)=1/(-2-h)

OpenStudy (zzr0ck3r):

what is the question?

OpenStudy (zzr0ck3r):

if \(h>0\) then \(f(h+2)=\dfrac{1}{|h+2|}=\dfrac{1}{h+2}\)

OpenStudy (nick5006):

oh whoops i was thinking of it differently, but i guess that's right.

OpenStudy (zzr0ck3r):

If we had \(h<-2\) then we would have \(x+2<0\) so \(f(h+2)=\dfrac{1}{h+2}=\dfrac{1}{-(h+2)}=\dfrac{1}{-h-2}=\dfrac{-1}{h+2} \)

OpenStudy (nick5006):

so that's the reason behind the answer?

OpenStudy (nick5006):

well i got it so thank you

OpenStudy (zzr0ck3r):

that was the reasoning behind a different answer.